# Fractional Sobolev paths on Wasserstein spaces and their energy-minimizing particle representations

### v1

@@ 2.7. Piecewise geodesic approximation on geodesic spaces
we define a continuous curve $X^{n}:[0,1]\to\mathcal{X}$ by
@@
and connecting by any constant-speed geodesic in between, i.e., the speed over each time segment is equal to $2^{n}d(x_{i},x_{i+1})$. In other words, $X^{n}$ is a piecewise geodesic connecting the points (2.55). In the next lemma, we compute $|X^{n}|_{b^{\alpha,p}}$.
@@ 2.8. Path measures and sample path regularity
from $(\mathcal{X}^{I},\mathcal{B}(\mathcal{X})^{I})\to(\bar{\mathbb{R}}_{+},\mathcal{B}(\bar{\mathbb{R}}_{+}))$ are not measurable. As functions from $(C(I;\mathcal{X}),\mathcal{C})\to(\bar{\mathbb{R}}_{+},\mathcal{B}(\bar{\mathbb{R}}_{+}))$, they are lower semi-continuous, in particular, measurable.
@@ 2.10. Tightness conditions for path measures
Let’s observe that a suitable Hölder regularity can guarantee the finiteness of the second part of (2.76). We thus arrive at the well-known Kolmogorov–Lamperti tightness result (cf. also [9, Corollary A.11]):
@@ 2.12. Compatible measures in $P_{p}(\mathcal{X})$
The compatibility property was implicitly defined in [7, Section 4] as an admissible property for a collection of transport maps and was further explored in [18]. What we outlined above can be generalized to more than 3 measures and to general spaces, for which we give the following definition, mentioned in [18, Remark 2.3.2]. Now, let $(\mathcal{X},d)$ be a complete separable metric space and $p\in[1,\infty)$.
@@
A prominent and well-known example of compatible measures is Wasserstein geodesics, as highlighted in the remark below. A compatible collection, however, need not lie on a Wasserstein geodesic. Take for instance an arbitrary measure and translate it along a non-geodesic curve. Another example is any collection of measures in $P_{p}(\mathbb{R})$ is compatible. As shown in [18, page 49], any collection of Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ that includes the standard Gaussian distribution and such that the covariance matrices of the measures are simultaneously diagonalizable (i.e. one can diagonalize all with one orthogonal matrix) is compatible. See more nontrivial examples presented in [7, Proposition 4.1] and explained in [18, Section 2.3.2].
@@ 3.1. Existence of a minimizer for the variational problem
Since the sublevels of $\Psi$ are compact (which follows from lower semi-continuity and relative compactness), the condition above implies that the family of measures $\{\pi_{n}\}_{n}\subset P(\Gamma_{T})$ is tight by Lemma 2.36. By Prokhorov’s theorem, there exists a subsequence $\{\pi_{n_{k}}\}_{k\in\mathbb{N}}$ such that $\pi_{n_{k}}\to\pi$ narrowly on $P(\Gamma_{T})$ as $k\to\infty$. By Lemma 2.43, we have $\pi\in\mathrm{Lift}(\mu_{t})$ and thus $\int_{\Gamma_{T}}\Psi{\mathrm{\,d}}\pi\geq M$. On the other hand, since $\Psi$ is lower semi-continuous, we obtain the reverse inequality
@@ 3.3. From Wasserstein curves to path measures: a superposition principle
Next, since $(\mathcal{X},d)$ is assumed to be a complete and locally compact length metric space, the Hopf–Rinow theorem 2.30 ensures the existence of geodesics between any two points. Therefore, we can consider the “geodesic interpolation” map
@@
and connecting by constant speed geodesics in between (i.e. the speed over each time segment is equal to $2^{n}d(x_{i},x_{i+1})$). In other words, $\ell_{\bm{x}}$ is a piecewise geodesic connecting the points $\bm{x}=(x_{0},\cdots,x_{N_{n}})$. Despite the possibility of non-uniqueness of geodesics, such a measurable geodesic selection map exists, as discussed in [3, Theorem 3.10 and Lemma 3.11]. Therefore, using this Borel map, we can build a sequence of path measures:
@@
Step 1 (Tightness of $\{\pi_{n}\}_{n\in\mathbb{N}}\subset P(\Gamma)$). We show that the family measures $\{\pi_{n}\}$ is tight in $P(C([0,1];\mathcal{X}))$. To this end, we use the tightness condition developed for path measures in Corollary 2.39 using the $b^{\alpha,p}$-semi-norm. Note that since $(\mathcal{X},d)$ is assumed to be a complete and locally compact length metric space, the Hopf–Rinow theorem 2.30 ensures that closed bounded sets in $\mathcal{X}$ are compact. Hence, we can use Corollary 2.39. Our goal is to show
@@ 3.6. Dynamic formulation of Wasserstein distance via Besov energy
As a consequence of Lemmas 3.11 and 3.12, we obtain the following generalization of the metric Benamou–Brenier formula to the fractional setting. Its proof is inspired by [2, Theorem 9.13].
@@
where $\ell:\mathcal{X}\times\mathcal{X}\to Geo([0,1];\mathcal{X})$ is the (Borel measurable) geodesic (selection) and interpolation map (as already used in this paper). Note that, by its very construction, $(e_{0},e_{1})_{\#}\pi^{*}\in\mathrm{OptCpl}(\mu,\nu)$, and $\pi^{*}$ is concentrated on $Geo([0,1];\mathcal{X})$. By Lemma 3.12 ($1\Rightarrow 2$), we have
@@ 4. Counterexamples
This elementary example is well-known in stochastic analysis, see e.g. [19, Chapter I. Exercise (2.12)], and it is studied in detail in [1, Example 1.1]. Consider the following collection of measures on $\mathcal{X}\coloneqq\mathbb{R}$:
@@
which clearly does not meet the Kolmogorov–Chentsov continuity criterion.
@@
Now take $p>1$ and note that the $p$-variation of the curve $(\mu_{t}^{j})$ with respect to $p$-Wasserstein metric is achieved by the partition $(\frac{i}{2^{j+1}})_{i\in\{0,\cdots 2^{j+1}\}}$. This can be verified using the interpolation result [9, Proposition 5.5]. We thus obtain
@@
Observation 4 ($W^{1,p}$-regularity of $(\mu_{t}^{j})$ remains constant). Consider $1<p<\infty$. Each $(\mu_{t}^{j})$ is an absolutely continuous curve in $p$-Wasserstein space with its metric speed $|\dot{\mu}_{t}^{j}|=1$ for almost all time for all $j$. Recalling [13, Theorem 5], we have here:

### v3

@@ 2.7. Piecewise geodesic approximation on geodesic spaces
let $X^{n}:[0,1]\to\mathcal{X}$ be a continuous curve that passes the points, i.e.,
@@
and connects in between by a constant-speed geodesic, i.e., the speed over each time segment is equal to $2^{n}d(x_{i},x_{i+1})$. In other words, $X^{n}$ is a piecewise geodesic connecting the points (2.55). In the next lemma, we compute $|X^{n}|_{b^{\alpha,p}}$. This quantity turns out to be independent of the choice of geodesics when geodesics are not unique.
@@ 2.8. Path measures and sample path regularity
from $(\mathcal{X}^{I},\mathcal{B}(\mathcal{X})^{I})\to(\bar{\mathbb{R}}_{+},\mathcal{B}(\bar{\mathbb{R}}_{+}))$ are not measurable. However, as functions from $(C(I;\mathcal{X}),\mathcal{C})\to(\bar{\mathbb{R}}_{+},\mathcal{B}(\bar{\mathbb{R}}_{+}))$, they are lower semi-continuous, in particular, measurable.
@@ 2.10. Tightness conditions for path measures
Let’s observe that a suitable Hölder regularity can guarantee the finiteness of the second part of (2.76). We thus arrive at the well-known Kolmogorov–Lamperti tightness result (cf. also [10, Corollary A.11]):
@@ 2.12. Compatible measures in $P_{p}(\mathcal{X})$
The compatibility property was implicitly defined in [7, Section 4] as an admissible property for a collection of transport maps and was further explored in [19]. What we outlined above can be generalized to more than 3 measures and to general spaces, for which we give the following definition, mentioned in [19, Remark 2.3.2]. Now, let $(\mathcal{X},d)$ be a complete separable metric space and $p\in[1,\infty)$.
@@
A prominent and well-known example of compatible measures is Wasserstein geodesics, as highlighted in the remark below. A compatible collection, however, need not lie on a Wasserstein geodesic. Take for instance an arbitrary measure and translate it along a non-geodesic curve. Another example is any collection of measures in $P_{p}(\mathbb{R})$ is compatible. As shown in [19, page 49], any collection of Gaussian measures on $\mathbb{R}^{\mathrm{d}}$ that includes the standard Gaussian distribution and such that the covariance matrices of the measures are simultaneously diagonalizable (i.e. one can diagonalize all with one orthogonal matrix) is compatible. See more nontrivial examples presented in [7, Proposition 4.1] and explained in [19, Section 2.3.2].
@@
2.13. Further remarks
In this paper, we frequently require a map that connects two points in a geodesic space via a geodesic. As in some geodesic spaces, there may be multiple geodesics connecting two points, we need a suitable selection map. The following remark is in this regard.
Remark 2.47 (Measurable selection of geodesics).
Let $(\mathcal{X},d)$ be a complete, separable, and geodesic metric space. Let $\Upsilon\in P(\mathcal{X}\times\mathcal{X})$. We frequently need to construct a path measure
$\pi\coloneqq(\ell)_{\#}\Upsilon\,\in P\big{(}(C([0,1];\mathcal{X}),\mathcal{C})\big{)}$ (2.92)
with a measurable geodesic selection map $\ell:\mathcal{X}\times\mathcal{X}\to Geo([0,1];\mathcal{X})$. We briefly emphasize that such a map exists and that the above is well-defined (using a similar argument as in the proof of [14, Proposition 1]). Define the multi-valued map $L:\mathcal{X}\times\mathcal{X}\to 2^{Geo([0,1];\mathcal{X})}$ as follows:
$L(x,y)\coloneqq\{\gamma\in Geo([0,1];\mathcal{X}):\quad\gamma_{0}=x,\gamma_{1}=y\}.$ (2.93)
This set is non-empty as $(\mathcal{X},d)$ is a geodesic space. The graph of $L$, denoted by
$\mathcal{G}(L)\coloneqq\Big{\{}(x,y,\gamma)\in\mathcal{X}\times\mathcal{X}\times C([0,1];\mathcal{X}):\quad\gamma\in L(x,y)\Big{\}},$ (2.94)
is a closed set. Indeed, we take a sequence $(x^{n},y^{n},\gamma^{n})_{n\in\mathbb{N}}\subset\mathcal{G}(L)$ convergent to $(x,y,\gamma)$ and show that $(x,y,\gamma)\in\mathcal{G}(L)$. First, we write
$d(x,\gamma_{0})$ $\leq d(x,x^{n})+d(x^{n},\gamma_{0})$ (2.95)
$=d(x,x^{n})+d(\gamma_{0}^{n},\gamma_{0}).$ (2.96)
Taking the limit $n\to\infty$, the second term on the right-hand side goes to zero as $d_{\infty}(\gamma^{n},\gamma)\to 0$. Thus, we have $\gamma_{0}=x$, and similarly $\gamma_{1}=y$. To show that $\gamma$ is a constant-speed geodesic, take any $t,s\in[0,1]$ and observe that
$d(\gamma_{s},\gamma_{t})$ $=\lim_{n\to\infty}d(\gamma^{n}_{s},\gamma^{n}_{t})$ (2.97)
$=\lim_{n\to\infty}|t-s|d(x^{n},y^{n})=|t-s|d(x,y).$ (2.98)
As $\mathcal{G}(L)$ is closed, it is Borel-measurable, i.e., it belongs to $\mathcal{B}(\mathcal{X}^{2}\times C)$, which is equal to $\mathcal{B}(\mathcal{X}^{2})\otimes\mathcal{C}$, since all spaces are separable here. Obviously, $\mathcal{G}(L)$ is also in $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}\otimes\mathcal{C}$, where $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}$ denotes the $\Upsilon$-completion of $\mathcal{B}(\mathcal{X}^{2})$. It follows from a measurable selection theorem by Aumann (see, e.g., [8, Theorem III.22] and [3, page 455]) that there exists a measurable map $\ell:\big{(}\mathcal{X}^{2},\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}\big{)}\to\big{(}C([0,1];\mathcal{X}),\mathcal{C}\big{)}$ such that $\ell(x,y)\in Geo([0,1];\mathcal{X})$ for $\Upsilon$-a.e. $(x,y)\in\mathcal{X}^{2}$. The above argument for selecting geodesics between two points can easily be generalized to multiple points.
@@ 3.1. Existence of a minimizer for the variational problem
Since the sublevels of $\Psi$ are compact (which follows from lower semi-continuity and relative compactness), the condition (3.2) implies that the family of measures $\{\pi_{n}\}_{n}\subset P(\Gamma_{T})$ is tight by Lemma 2.36. By Prokhorov’s theorem, there exists a subsequence $\{\pi_{n_{k}}\}_{k\in\mathbb{N}}$ such that $\pi_{n_{k}}\to\pi$ narrowly on $P(\Gamma_{T})$ as $k\to\infty$. By Lemma 2.43, we have $\pi\in\mathrm{Lift}(\mu_{t})$ and thus $\int_{\Gamma_{T}}\Psi{\mathrm{\,d}}\pi\geq M$. On the other hand, since $\Psi$ is lower semi-continuous, we obtain the reverse inequality
@@ 3.3. From Wasserstein curves to path measures: a superposition principle
Next, since $(\mathcal{X},d)$ is assumed to be a complete and locally compact length metric space, the Hopf–Rinow Theorem 2.30 ensures the existence of at least one geodesic between any two points. We consider a geodesic selection and interpolation map connecting the points with constant-speed geodesics:
@@
and connecting by a constant speed geodesics in between (i.e. the speed over each time segment is equal to $2^{n}d(x_{i},x_{i+1})$). In other words, $\ell_{\bm{x}}$ is a piecewise geodesic connecting the points. As discussed in Remark 2.47 for the two-point case, one can always find a $(\mathcal{B}(\mathcal{X}^{2^{n}+1})_{\Upsilon_{n}},\mathcal{C})$-measurable geodesic selection map $\ell_{\bm{x}}$, despite the possible non-uniqueness of geodesics. Here, $\mathcal{B}(\mathcal{X}^{2^{n}+1})_{\Upsilon_{n}}$ denotes the $\Upsilon_{n}$-completion of $\mathcal{B}(\mathcal{X}^{2^{n}+1})$. Therefore, using this measurable map, we can build a sequence of path measures:
@@
Step 1 (Tightness of $\{\pi_{n}\}_{n\in\mathbb{N}}\subset P(\Gamma)$). We show that the family measures $\{\pi_{n}\}$ is tight in $P(C([0,1];\mathcal{X}))$. To this end, we use the tightness condition developed for path measures in Corollary 2.39 using the $b^{\alpha,p}$-norm. Note that since $(\mathcal{X},d)$ is assumed to be a complete and locally compact length metric space, the Hopf–Rinow theorem 2.30 ensures that closed bounded sets in $\mathcal{X}$ are compact. Hence, we can use Corollary 2.39. Our goal is to show
@@ 3.6. Dynamic formulation of Wasserstein distance via Besov energy
As a consequence of Lemmas 3.11 and 3.12, we obtain a generalization of the metric Benamou–Brenier formula to the fractional setting. Its proof is similar to that for $W^{1,p}$-energy (e.g. [2, Theorem 9.13]).
@@
where $\ell:\mathcal{X}\times\mathcal{X}\to Geo([0,1];\mathcal{X})$ is a $(\mathcal{B}(\mathcal{X}^{2})_{\Upsilon},\mathcal{C})$-measurable geodesic selection, whose existence is discussed in Remark 2.47. Here, $\mathcal{B}(\mathcal{X}^{2})_{\Upsilon}$ denotes the $\Upsilon$-completion of $\mathcal{B}(\mathcal{X}^{2})$.
Note that, by its very construction, $(e_{0},e_{1})_{\#}\pi^{*}\in\mathrm{OptCpl}(\mu,\nu)$, and $\pi^{*}$ is concentrated on $Geo([0,1];\mathcal{X})$. By Lemma 3.12 ($1\Rightarrow 2$), we have
@@ 4. Counterexamples
This elementary example is well-known in stochastic analysis, see e.g. [20, Chapter I. Exercise (2.12)], and it is studied in detail in [1, Example 1.1]. Consider the following collection of measures on $\mathcal{X}\coloneqq\mathbb{R}$:
@@
which clearly does not meet the Kolmogorov–Čentsov continuity criterion.
@@
Now take $p>1$ and note that the $p$-variation of the curve $(\mu_{t}^{j})$ with respect to $p$-Wasserstein metric is achieved by the partition $(\frac{i}{2^{j+1}})_{i\in\{0,\cdots 2^{j+1}\}}$. This can be verified using the interpolation result [10, Proposition 5.5]. We thus obtain
@@
Observation 4 ($W^{1,p}$-regularity of $(\mu_{t}^{j})$ remains constant). Consider $1<p<\infty$. Each $(\mu_{t}^{j})$ is an absolutely continuous curve in $p$-Wasserstein space with its metric speed $|\dot{\mu}_{t}^{j}|=1$ for almost all time for all $j$. Recalling [14, Theorem 5], we have here:
@@
Data availability. Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.